# Supersymmetry and Localization on Three-Dimensional Orbifolds

### v1

@@ 5.1.1 General Construction
In ${\left({\Theta,\sigma,y}\right)}$ coordinates the line element of $S^{3}_{q,p}$ reads
${\rm d}s^{\prime 2}$ $=c_{11}{\rm d}\Theta^{2}+c_{22}{\rm d}\sigma^{2}+h{\left({{\rm d}y+A}\right)}^{2}~{},$
$c_{11}$ $=\frac{1}{4}{\left({q^{2}\cos^{2}\frac{\Theta}{2}+p^{2}\sin^{2}\frac{\Theta}{2}}\right)}=f^{2}/4~{},$
$c_{22}$ $=\frac{p^{2}q^{2}\sin^{2}\Theta}{4{\left({p^{2}t_{22}^{2}\cos^{2}\frac{\Theta}{2}+q^{2}t_{12}^{2}\sin^{2}\frac{\Theta}{2}}\right)}}~{},$
$h$ $={\left({p^{2}t_{22}^{2}\cos^{2}\frac{\Theta}{2}+q^{2}t_{12}^{2}\sin^{2}\frac{\Theta}{2}}\right)}~{},$
$A$ $=\frac{p^{2}t_{21}t_{22}\cos^{2}\frac{\Theta}{2}+q^{2}t_{11}t_{12}\sin^{2}\frac{\Theta}{2}}{p^{2}t_{22}^{2}\cos^{2}\frac{\Theta}{2}+q^{2}t_{12}^{2}\sin^{2}\frac{\Theta}{2}}{\rm d}\sigma~{}.$ (5.4)
Near $\Theta=0$ the metric becomes
@@
${\rm d}s^{2}_{N}$ $=\frac{q^{2}}{4}{\left({{\rm d}\Theta^{2}+\frac{\Theta^{2}}{t_{22}^{2}}{\rm d}\sigma^{2}}\right)}+p^{2}t_{22}^{2}{\left({{\rm d}\varphi+\frac{t_{21}}{t_{22}}{\rm d}\sigma}\right)}^{2}~{},$ (5.5)
while at $\Theta=\pi$ the metric reads
@@
${\rm d}s^{2}_{S}$ $=\frac{p^{2}}{4}{\left({{\rm d}\Theta^{2}+\frac{{\left({\pi-\Theta}\right)}^{2}}{t_{12}^{2}}{\rm d}\sigma^{2}}\right)}+q^{2}t_{12}^{2}{\left({{\rm d}y+\frac{t_{11}}{t_{12}}{\rm d}\sigma}\right)}^{2}~{},$ (5.6)
@@
$t_{12}=n_{S}~{},$ $t_{22}=n_{N}~{}.$ (5.7)
The orbifold Euler-characteristic of the base is
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}\Theta{\rm d}z\sqrt{g}R=\frac{1}{|t_{12}|}+\frac{1}{|t_{22}|}=\frac{1}{n_{S}}+\frac{1}{n_{N}}=\chi_{\rm orb}~{}.$ (5.8)
Moreover,
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}A=\frac{t_{11}}{t_{12}}-\frac{t_{21}}{t_{22}}=\frac{\det t}{t_{12}t_{22}}=\frac{1}{n_{N}n_{S}}~{},$ (5.9)
where we used
$t_{11}t_{22}-t_{21}t_{12}=n_{N}t_{S}-n_{S}t_{N}=1~{},$ (5.10)
with the relabelling
$t_{11}=t_{S}~{},$ $t_{21}=t_{N}~{},$ (5.11)
giving
$t=\begin{pmatrix}t_{S}&n_{S}\\ t_{N}&n_{N}\end{pmatrix}~{},$ $\det t=n_{N}t_{S}-n_{S}t_{N}=1~{}.$ (5.14)
In ${\left({\sigma,y}\right)}$ coordinates, the phase of the Killing spinor $\zeta$ becomes
$\exp{\left({\frac{{\rm i}}{2}\alpha_{2}\phi_{1}+\frac{{\rm i}}{2}\alpha_{3}\phi_{2}}\right)}$ $=\exp{\left[{\frac{{\rm i}}{2}{\left({t_{11}\alpha_{2}+t_{21}\alpha_{3}}\right)}\sigma+\frac{{\rm i}}{2}{\left({t_{12}\alpha_{2}+t_{22}\alpha_{3}}\right)}y}\right]}~{}.$ (5.15)
We introduce
$2c_{R}$ $=t_{11}\alpha_{2}+t_{21}\alpha_{3}~{},$
$2\lambda_{R}$ $=t_{12}\alpha_{2}+t_{22}\alpha_{3}~{},$ (5.16)
with inverse relations
$\alpha_{2}$ $=2{\left({t_{22}c_{R}-t_{21}\lambda_{R}}\right)}~{},$
$\alpha_{3}$ $=2{\left({-t_{12}c_{R}+t_{11}\lambda_{R}}\right)}~{}.$ (5.17)
Invariance under $y\sim y+2\pi$ yields
$t_{12}\alpha_{2}+t_{22}\alpha_{3}=2\lambda_{R}~{},$ $\lambda_{R}\in\mathbb{Z}~{}.$ (5.18)
The background field $A^{C}$ is smooth on $S^{3}_{q,p}$ iff $\alpha_{2}=-1$ and $\alpha_{3}=-\mathfrak{s}_{w}$, hence
$-n_{S}-\mathfrak{s}_{w}n_{N}=2\lambda_{R}~{},$ $\mathfrak{s}_{w}=\pm 1~{}.$ (5.19)
If $\gcd{\left({n_{N},n_{S}}\right)}=1$, then both $n_{N}$ and $n_{S}$ are odd. If we allow for spinors that are anti-periodic along $y$, then either $n_{N}$ or $n_{S}$ can be even. If we denote by $c_{R}$ the coefficient of $\sigma$ in the phase of the Killing spinor $\zeta$, thus
$2c_{R}$ $=t_{11}\alpha_{2}+t_{21}\alpha_{3}=-t_{S}-\mathfrak{s}_{w}t_{N}~{},$
$2\lambda_{R}$ $=t_{12}\alpha_{2}+t_{22}\alpha_{3}=-n_{S}-\mathfrak{s}_{w}n_{N}~{},$ (5.20)
with inverse relations
$\alpha_{2}$ $=2{\left({t_{22}c_{R}-t_{21}\lambda_{R}}\right)}=2{\left({n_{N}c_{R}-t_{N}\lambda_{R}}\right)}~{},$
$\alpha_{3}$ $=2{\left({-t_{12}c_{R}+t_{11}\lambda_{R}}\right)}=2{\left({-n_{S}c_{R}+t_{S}\lambda_{R}}\right)}~{}.$ (5.21)
In ${\left({\sigma,y}\right)}$ coordinates, the phase of the Fourier modes on $S^{3}_{q,p}$ reads
$\exp{\left({{\rm i}m_{i}\phi_{1}+{\rm i}n_{i}\phi_{2}}\right)}$ $=\exp{\left[{{\rm i}{\left({t_{11}m_{i}+t_{21}n_{i}}\right)}\sigma+{\rm i}{\left({t_{12}m_{i}+t_{22}n_{i}}\right)}y}\right]}=\exp{\left[{{\rm i}j_{i}\sigma+{\rm i}\mathfrak{m}_{i}y}\right]}~{},$ (5.22)
with
$j_{i}$ $=t_{11}m_{i}+t_{21}n_{i}=t_{S}m_{i}+t_{N}n_{i}\in\mathbb{Z}~{},$
$\mathfrak{m}_{i}$ $=t_{12}m_{i}+t_{22}n_{i}=n_{S}m_{i}+n_{N}n_{i}\in\mathbb{Z}~{},$ (5.23)
and inverse relations
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},$
$n_{i}$ $=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{}.$ (5.24)
The Killing vector in ${\left({\sigma,y}\right)}$ coordinates is
@@
$K$ $=\mathcal{N}p{\left({\frac{\mathfrak{s}_{w}}{q}\partial_{\phi_{1}}+\frac{1}{p}\partial_{\phi_{2}}}\right)}~{},$
$=\mathcal{N}p{\left[{\frac{\mathfrak{s}_{w}}{q}{\left({t_{22}\partial_{\sigma}-t_{21}\partial_{y}}\right)}+\frac{1}{p}{\left({-t_{12}\partial_{\sigma}+t_{11}\partial_{y}}\right)}}\right]}~{},$
$=\mathcal{N}p{\left[{{\left({\frac{\mathfrak{s}_{w}t_{22}}{q}-\frac{t_{12}}{p}}\right)}\partial_{\sigma}+{\left({\frac{t_{11}}{p}-\frac{\mathfrak{s}_{w}t_{21}}{q}}\right)}\partial_{y}}\right]}~{},$
$=\mathcal{N}p{\left({\frac{t_{11}}{p}-\frac{\mathfrak{s}_{w}t_{21}}{q}}\right)}{\left({{\rm i}\omega_{\rm 2d}\partial_{\sigma}+\partial_{y}}\right)}~{},$ (5.25)
where we defined the angular-momentum fugacity on the base $\mathbb{\Sigma}$ as
$\alpha_{\rm 2d}={\rm i}\omega_{\rm 2d}=\frac{\mathfrak{s}_{w}t_{22}p-t_{12}q}{t_{11}q-\mathfrak{s}_{w}t_{21}p}=\frac{\mathfrak{s}_{w}n_{N}p-n_{S}q}{t_{S}q-\mathfrak{s}_{w}t_{N}p}~{}.$ (5.26)
In the new variables the one-loop determinant becomes
$Z_{\text{1-L}}$ $=\frac{\prod_{\ell_{2},k_{2}\in\mathbb{N}}{\left({p\ell_{2}+qk_{2}-{\rm i}\mathfrak{s}_{w}pq\sigma_{c}-\frac{q_{R}^{B}}{2}{\left({p+q}\right)}}\right)}}{\prod_{\ell_{1},k_{1}\in\mathbb{N}}{\left({p\ell_{1}+qk_{1}+{\rm i}\mathfrak{s}_{w}pq\sigma_{c}+\frac{q_{R}^{\phi}}{2}{\left({p+q}\right)}}\right)}}~{},$
@@
$=\frac{\prod_{m_{2},n_{2}}{\left[{-pm_{2}-\mathfrak{s}_{w}qn_{2}-{\rm i}\mathfrak{s}_{w}pq\sigma_{c}+\frac{q_{R}^{B}}{2}{\left({p\alpha_{2}+q\mathfrak{s}_{w}\alpha_{3}}\right)}}\right]}}{\prod_{m_{1},n_{1}}{\left[{pm_{1}+\mathfrak{s}_{w}qn_{1}+{\rm i}\mathfrak{s}_{w}pq\sigma_{c}-\frac{q_{R}^{\phi}}{2}{\left({p\alpha_{2}+q\mathfrak{s}_{w}\alpha_{3}}\right)}}\right]}}~{},$

### v2

@@ 5.1 General construction
In ${\left({\vartheta,\varphi,\psi}\right)}$ coordinates the line element of $S^{3}_{b_{s},b_{c}}$ reads
${\rm d}s^{\prime 2}$ $={\left({f^{2}/4}\right)}{\rm d}\vartheta^{2}+{\rm h}_{11}{\rm d}\varphi^{2}+{\rm h}_{22}{\left({{\rm d}\psi+A^{(1)}}\right)}^{2}~{},$
${\rm h}_{11}$ $=\frac{b_{s}^{2}b_{c}^{2}\,\sin^{2}\vartheta}{4{\left({b_{s}^{2}\,n_{1}^{2}\cos^{2}\frac{\vartheta}{2}+b_{c}^{2}\,n_{2}^{2}\sin^{2}\frac{\vartheta}{2}}\right)}}~{},$
${\rm h}_{22}$ $=b_{s}^{2}\,n_{1}^{2}\cos^{2}\frac{\vartheta}{2}+b_{c}^{2}\,n_{2}^{2}\sin^{2}\frac{\vartheta}{2}~{},$
$A^{(1)}$ $=\frac{b_{s}^{2}\,n_{1}t_{1}\cos^{2}\frac{\vartheta}{2}+b_{c}^{2}\,n_{2}t_{2}\sin^{2}\frac{\vartheta}{2}}{b_{s}^{2}\,n_{1}^{2}\cos^{2}\frac{\vartheta}{2}+b_{c}^{2}\,n_{2}^{2}\sin^{2}\frac{\vartheta}{2}}{\rm d}\varphi~{}.$ (5.6)
Near $\vartheta=0$ the metric becomes
${\rm d}s^{2}|_{\mathcal{U}_{+}}$ $=\frac{b_{c}^{2}}{4}{\left({{\rm d}\vartheta^{2}+\frac{\vartheta^{2}}{n_{1}^{2}}{\rm d}\varphi^{2}}\right)}+b_{s}^{2}\,n_{1}^{2}{\left({{\rm d}\psi+\frac{t_{1}}{n_{1}}{\rm d}\varphi}\right)}^{2}~{},$ (5.7)
while at $\vartheta=\pi$ the metric reads
${\rm d}s^{2}|_{\mathcal{U}_{-}}$ $=\frac{b_{s}^{2}}{4}{\left({{\rm d}\vartheta^{2}+\frac{{\left({\pi-\vartheta}\right)}^{2}}{n_{2}^{2}}{\rm d}\varphi^{2}}\right)}+b_{c}^{2}\,n_{2}^{2}{\left({{\rm d}\psi+\frac{t_{2}}{n_{2}}{\rm d}\varphi}\right)}^{2}~{},$ (5.8)
corresponding to a $\mathcal{O}{\left({-1}\right)}$ circle bundle whose base has the topology of a spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{n_{1},n_{2}}\right]}}$: indeed, the orbifold Euler characteristic of the base is
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}\vartheta{\rm d}\varphi\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=\frac{1}{n_{2}}+\frac{1}{n_{1}}~{},$ (5.9)
while the flux of $A^{(1)}$ reads
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}A^{(1)}=\frac{t_{2}}{n_{2}}-\frac{t_{1}}{n_{1}}=\frac{1}{n_{1}n_{2}}~{}.$ (5.10)
In terms of ${\left({\varphi,\psi}\right)}$ the complex coordinates (5.3) read
$z_{1}=b_{s}e^{{\rm i}{\left({n_{1}\psi+t_{1}\varphi}\right)}}\sin x~{},\qquad z_{2}=b_{c}e^{{\rm i}{\left({n_{2}\psi+t_{2}\varphi}\right)}}\cos x~{},$ (5.11)
which, under a shift along the fiber $\psi\to\psi+\Delta\psi$ with $\Delta\psi\in\mathbb{R}$ and ${\left\llbracket{\Delta\psi}\right\rrbracket_{{\left({2\pi}\right)}}}\neq 0$, behave as
$z_{1}\to\lambda_{\psi}^{n_{1}}z_{1}~{},\qquad z_{2}\to\lambda_{\psi}^{n_{2}}z_{2}~{},$ $\lambda_{\psi}=e^{{\rm i}\,\Delta\psi}\in U(1)\subset\mathbb{C}^{*}~{}.$ (5.12)
The base $\mathbb{\Sigma}$ is given by the quotient of $S^{3}_{b_{s},b_{c}}\subset\mathbb{C}^{2}$ with respect to the $\mathbb{C}^{*}$ action generated by $\lambda_{\psi}$, where the coordinates (5.3) satisfy the identification
${\left({z_{1},z_{2}}\right)}\sim{\left({\lambda_{\psi}^{n_{1}}z_{1},\lambda_{\psi}^{n_{2}}z_{2}}\right)}~{},$ (5.13)
again confirming that $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{n_{1},n_{2}}\right]}}$. The spinors
$\zeta_{\alpha}=e^{\frac{{\rm i}}{2}\alpha_{2}\varphi_{1}+\frac{{\rm i}}{2}\alpha_{3}\varphi_{2}}\begin{pmatrix}\sqrt{1+\sin x}\\ \frac{\cos x}{\sqrt{1+\sin x}}\end{pmatrix}_{\alpha}~{},$ $\widetilde{\zeta}_{\alpha}={\mathfrak{s}_{\omega}}e^{-\frac{{\rm i}}{2}\alpha_{2}\varphi_{1}-\frac{{\rm i}}{2}\alpha_{3}\varphi_{2}}\begin{pmatrix}\frac{\cos x}{\sqrt{1+\sin x}}\\ -\sqrt{1+\sin x}\end{pmatrix}_{\alpha}~{},$ (5.14)
with ${\mathfrak{s}_{\omega}}=\pm 1$, fulfil the Killing spinor equation on $S^{3}_{b_{s},b_{c}}$ with background fields
$A^{C}$ $=\frac{1}{2}{\left({\frac{b_{s}}{f}+\alpha_{2}}\right)}{\rm d}\varphi_{1}+\frac{1}{2}{\left({\frac{{\mathfrak{s}_{\omega}}b_{c}}{f}+\alpha_{3}}\right)}{\rm d}\varphi_{2}~{},$
$V$ $=-\frac{b_{s}}{f}{\left({1+{\rm i}\,{\mathfrak{s}_{\omega}}fH}\right)}\sin^{2}x\,{\rm d}\varphi_{1}-\frac{b_{c}}{f}{\left({{\mathfrak{s}_{\omega}}+{\rm i}fH}\right)}\cos^{2}x\,{\rm d}\varphi_{2}~{},$ (5.15)
and $A=A^{C}+(3\,V/2)$. Killing spinors and background fields are smooth on the whole $S^{3}_{b_{s},b_{c}}$ if $\alpha_{2}=1$ and $\alpha_{3}={\mathfrak{s}_{\omega}}$. In ${\left({\varphi,\psi}\right)}$ coordinates the phase of the Killing spinors reads
$\exp{\left({\frac{{\rm i}}{2}\alpha_{2}\varphi_{1}+\frac{{\rm i}}{2}\alpha_{3}\varphi_{2}}\right)}$ $=\exp{\left[{\frac{{\rm i}}{2}{\left({t_{1}\alpha_{2}+t_{2}\alpha_{3}}\right)}\varphi+\frac{{\rm i}}{2}{\left({n_{1}\alpha_{2}+n_{2}\alpha_{3}}\right)}\psi}\right]}~{}.$ (5.16)
If we denote by $c_{R}$ and $\lambda_{R}$ the coefficient of $\varphi$ and $\psi$ in the phase above, we have
$c_{R}$ $=\frac{1}{2}{\left({t_{1}\alpha_{2}+t_{2}\alpha_{3}}\right)}=\frac{1}{2}{\left({t_{1}+{\mathfrak{s}_{\omega}}t_{2}}\right)}~{},$
$\lambda_{R}$ $=\frac{1}{2}{\left({n_{1}\alpha_{2}+n_{2}\alpha_{3}}\right)}=\frac{1}{2}{\left({n_{1}+{\mathfrak{s}_{\omega}}n_{2}}\right)}~{},$ (5.17)
with inverse relations
$-2n_{2}c_{R}+2t_{2}\lambda_{R}$ $=\alpha_{2}=1~{},$
$2n_{1}c_{R}-2t_{1}\lambda_{R}$ $=\alpha_{3}={\mathfrak{s}_{\omega}}~{}.$ (5.18)
In particular, $\lambda_{R}$ is an $R$-symmetry flat connection along the $S^{1}$ fiber parameterized. In [34] flat connections along the Hopf fiber were shown to be related to fluxes on the base. In fact, $\lambda_{R}$ precisely coincides with ${\left({n_{1}n_{2}}\right)}$ times the $R$-symmetry flux affecting the Killing spinors on the spindle base, where ${\mathfrak{s}_{\omega}}=\pm 1$ correspond to twist and anti-twist, respectively.
The eigenfunctions of $\delta^{2}=K^{\mu}D_{\mu}$ contributing to the one-loop determinant of a chiral multiplet of $R$-charge $r$ on $S^{3}_{b_{s},b_{c}}$ have the form
$\mathcal{B}_{m_{1}^{(1)},m_{2}^{(1)}}{\left({x,\varphi_{1},\varphi_{2}}\right)}$ $=e^{{\rm i}m_{1}^{(1)}\varphi_{1}+{\rm i}m_{2}^{(1)}\varphi_{2}}B_{m_{1}^{(1)},m_{2}^{(1)}}{\left({x}\right)}~{},$
$\Phi_{m_{1}^{(2)},m_{2}^{(2)}}{\left({x,\varphi_{1},\varphi_{2}}\right)}$ $=e^{{\rm i}m_{1}^{(2)}\varphi_{1}+{\rm i}m_{2}^{(2)}\varphi_{2}}\phi_{m_{1}^{(2)},m_{2}^{(2)}}{\left({x}\right)}~{},$ (5.19)
where the modes $\mathcal{B}_{m_{1}^{(1)},m_{2}^{(1)}}$ have $R$-charge ${\left({r-2}\right)}$ while $\Phi_{m_{1}^{(2)},m_{2}^{(2)}}$ have $R$-charge $r$. The conditions $\mathcal{B}_{m_{1}^{(1)},m_{2}^{(1)}}\in{\rm Ker}\,L_{P}$ and $\Phi_{m_{1}^{(2)},m_{2}^{(2)}}\in{\rm Ker}\,L_{\widetilde{P}}$ require
$m_{1}^{(1)}=k_{1}~{},$ $m_{2}^{(1)}={\mathfrak{s}_{\omega}}k_{2}~{},$
$m_{1}^{(2)}=-\ell_{1}~{},$ $m_{2}^{(2)}=-{\mathfrak{s}_{\omega}}\ell_{2}~{},$ (5.20)
with $k_{1},k_{2},\ell_{1},\ell_{2}\in\mathbb{N}$. In ${\left({\varphi,\psi}\right)}$ coordinates the phases of the Fourier modes of the cohomological fields $\mathcal{B}_{m_{1}^{(1)},m_{2}^{(1)}}$ and $\Phi_{m_{1}^{(2)},m_{2}^{(2)}}$ read
$\exp{\left({{\rm i}\,m_{1}^{(i)}\varphi_{1}+{\rm i}\,m_{2}^{(i)}\varphi_{2}}\right)}=\exp{\left({{\rm i}\,j^{(i)}\varphi+{\rm i}\,\mathfrak{f}^{(i)}\psi}\right)}~{},$ (5.21)
with $i=1,2$ and
$j^{(i)}=t_{1}m_{1}^{(i)}+t_{2}m_{2}^{(i)}~{},$ $\mathfrak{f}^{(i)}=n_{1}m_{1}^{(i)}+n_{2}m_{2}^{(i)}~{},$ (5.22)
along with the inverse relations
$m_{1}^{(i)}=-n_{2}j^{(i)}+t_{2}\mathfrak{f}^{(i)}~{},$ $m_{2}^{(i)}=n_{1}j^{(i)}-t_{1}\mathfrak{f}^{(i)}~{}.$ (5.23)
In analogy with $\lambda_{R}$, we expect that $\mathfrak{f}^{(1)}$ and $\mathfrak{f}^{(2)}$ corresponds to ${\left({n_{1}n_{2}}\right)}$ times the flux affecting $\mathcal{B}_{m_{1}^{(1)},m_{2}^{(1)}}$ and $\Phi_{m_{1}^{(2)},m_{2}^{(2)}}$, respectively. Similarly, the Killing vector $K=(\zeta\sigma^{\mu}\widetilde{\zeta})\partial_{\mu}$ written in ${\left({\varphi,\psi}\right)}$ coordinates is
@@
$K$ $=k_{0}b_{c}{\left({\frac{{\mathfrak{s}_{\omega}}}{b_{s}}\partial_{\varphi_{1}}+\frac{1}{b_{c}}\partial_{\varphi_{2}}}\right)}=k_{0}^{\prime}{\left({{\omega}_{\rm 2d}\partial_{\varphi}+\partial_{\psi}}\right)}~{},$ (5.24)
where we introduced
$k_{0}^{\prime}=\frac{k_{0}{\left({t_{2}{\mathfrak{s}_{\omega}}b_{c}-t_{1}b_{s}}\right)}}{b_{s}}~{},$ ${\omega}_{\rm 2d}=\frac{n_{1}b_{s}-{\mathfrak{s}_{\omega}}n_{2}b_{c}}{{\mathfrak{s}_{\omega}}t_{2}b_{c}-t_{1}b_{s}}~{}.$ (5.25)
The one-loop determinant of a chiral multiplet coupled to an Abelian vector multiplet on $S^{3}_{b_{s},b_{c}}$ is
@@
$Z_{\text{1-L}}^{\rm CM}$ $=\frac{\prod_{k_{1},k_{2}\in\mathbb{N}}{\left[{b_{c}k_{1}+b_{s}k_{2}+{\left({1-\frac{r}{2}}\right)}{\left({b_{c}+b_{s}}\right)}-{\rm i}{\mathfrak{s}_{\omega}}b_{c}b_{s}\sigma_{c}}\right]}}{\prod_{\ell_{1},\ell_{2}\in\mathbb{N}}{\left[{b_{c}\ell_{1}+b_{s}\ell_{2}+\frac{r}{2}{\left({b_{c}+b_{s}}\right)}+{\rm i}{\mathfrak{s}_{\omega}}b_{c}b_{s}\sigma_{c}}\right]}}~{},$
$=\frac{\prod_{m^{(1)}_{1},m^{(1)}_{2}}{\left[{b_{c}m^{(1)}_{1}+{\mathfrak{s}_{\omega}}b_{s}m^{(1)}_{2}+{\left({1-\frac{r}{2}}\right)}{\left({b_{c}\alpha_{2}+{\mathfrak{s}_{\omega}}b_{s}\alpha_{3}}\right)}-{\rm i}{\mathfrak{s}_{\omega}}b_{c}b_{s}\sigma_{c}}\right]}}{\prod_{m^{(2)}_{1},m^{(2)}_{2}}{\left[{-b_{c}m^{(2)}_{1}-{\mathfrak{s}_{\omega}}b_{s}m^{(2)}_{2}+\frac{r}{2}{\left({b_{c}\alpha_{2}+{\mathfrak{s}_{\omega}}b_{s}\alpha_{3}}\right)}+{\rm i}{\mathfrak{s}_{\omega}}b_{c}b_{s}\sigma_{c}}\right]}}~{},$ (5.26)
which, by means of (5.1) and (5.23), can be reformulated exclusively in terms of two-dimensional quantities:
$Z_{\rm 2d}^{\rm CM}$ $=\frac{\prod_{j^{(1)}}{\left\{{{\omega}_{\rm 2d}{\left[{j^{(1)}+{\left({2-r}\right)}c_{R}}\right]}+\mathfrak{f}^{(1)}+{\left({2-r}\right)}\lambda_{R}+{\rm i}\,\sigma_{\rm 2d}}\right\}}}{\prod_{j^{(2)}}{\left[{{\omega}_{\rm 2d}{\left({j^{(2)}-r\,c_{R}}\right)}+\mathfrak{f}^{(2)}-r\lambda_{R}+{\rm i}\,\sigma_{\rm 2d}}\right]}}~{},$ (5.27)
with
$\sigma_{\rm 2d}=\frac{b_{c}b_{s}\sigma_{c}}{t_{1}b_{s}-{\mathfrak{s}_{\omega}}t_{2}b_{c}}~{},$ (5.28)
while the ranges of $j^{(1)}$ and $j^{(2)}$ are determined by
$-n_{2}j^{(1)}+t_{2}\mathfrak{f}^{(1)}=k_{1}~{},$ $n_{1}j^{(1)}-t_{1}\mathfrak{f}^{(1)}={\mathfrak{s}_{\omega}}k_{2}~{},$
$-n_{2}j^{(2)}+t_{2}\mathfrak{f}^{(2)}=-\ell_{1}~{},$ $-n_{1}j^{(2)}+t_{1}\mathfrak{f}^{(2)}={\mathfrak{s}_{\omega}}\ell_{2}~{},$ (5.29)
taking into account that $k_{1},k_{2},\ell_{1}$ and $\ell_{2}$ are all integers. We recall that the numerator of $Z_{\text{1-L}}^{\rm CM}$ arises from the fermionic modes associated to the chiral-multiplet cohomological field $B$ defined in Appendix E, while the denominator corresponds to the contribution of bosonic degrees of freedom encoded by the chiral-multiplet scalar $\phi$. This is consistent with relating the integers ${\left({j^{(1)},\mathfrak{f}^{(1)}}\right)}$ and ${\left({j^{(2)},\mathfrak{f}^{(2)}}\right)}$ with $B$ and $\phi$, respectively. We now proceed to apply the formula (5.27) to the cases where the bases of the fibrations are either topologically twisted or anti-twisted spindles.
5.2 Topologically twisted spindle
If ${\mathfrak{s}_{\omega}}=+1$, then
$c_{R}=\frac{1}{2}{\left({t_{1}+t_{2}}\right)}~{},$ $\lambda_{R}=\frac{1}{2}{\left({n_{1}+n_{2}}\right)}~{},$ (5.30)
and
$-n_{2}j^{(1)}+t_{2}\mathfrak{f}^{(1)}\geq 0~{},$ $n_{1}j^{(1)}-t_{1}\mathfrak{f}^{(1)}\geq 0~{},$
$-n_{2}j^{(2)}+t_{2}\mathfrak{f}^{(2)}\leq 0~{},$ $-n_{1}j^{(2)}+t_{1}\mathfrak{f}^{(2)}\geq 0~{},$ (5.31)
@@ 5.2 Topologically twisted spindle
${\left\lceil{t_{1}\mathfrak{f}^{(1)}/n_{1}}\right\rceil}\leq j^{(1)}\leq{\left\lfloor{t_{2}\mathfrak{f}^{(1)}/n_{2}}\right\rfloor}~{},$ ${\left\lceil{t_{2}\mathfrak{f}^{(2)}/n_{2}}\right\rceil}\leq j^{(2)}\leq{\left\lfloor{t_{1}\mathfrak{f}^{(2)}/n_{1}}\right\rfloor}~{}.$ (5.32)
As anticipated, we interpret $\mathfrak{f}^{(1)}$ and $\mathfrak{f}^{(2)}$ as ${\left({n_{1}n_{2}}\right)}$ times the total fluxes attached to $B$ and $\phi$:
$\mathfrak{f}^{(1)}={\left({r-2}\right)}\lambda_{R}+\mathfrak{m}~{},$ $\mathfrak{f}^{(2)}=r\,\lambda_{R}+\mathfrak{m}~{},$ (5.33)
with $\mathfrak{m}\in\mathbb{Z}$. Thanks to the identities
@@
${\left\lceil{\frac{t_{1}\mathfrak{f}^{(1)}}{n_{1}}}\right\rceil}$ $=1+{\left\lfloor{\frac{\mathfrak{p}_{1}}{n_{1}}}\right\rfloor}+{\left({r-2}\right)}c_{R}~{},$
${\left\lfloor{\frac{t_{2}\mathfrak{f}^{(1)}}{n_{2}}}\right\rfloor}$ $=-1-{\left\lfloor{-\frac{\mathfrak{p}_{2}}{n_{2}}}\right\rfloor}+{\left({r-2}\right)}c_{R}~{},$
${\left\lceil{\frac{t_{2}\mathfrak{f}^{(2)}}{n_{2}}}\right\rceil}$ $={\left\lceil{\frac{\mathfrak{p}_{2}}{n_{2}}}\right\rceil}+r\,c_{R}~{},$
${\left\lfloor{\frac{t_{1}\mathfrak{f}^{(2)}}{n_{1}}}\right\rfloor}$ $={\left\lfloor{\frac{\mathfrak{p}_{1}}{n_{1}}}\right\rfloor}+r\,c_{R}~{},$ (5.34)
we obtain
$1+{\left\lfloor{\frac{\mathfrak{p}_{1}}{n_{1}}}\right\rfloor}\leq\ell^{(1)}\leq-1-{\left\lfloor{-\frac{\mathfrak{p}_{2}}{n_{2}}}\right\rfloor}~{},$ $-{\left\lfloor{-\frac{\mathfrak{p}_{2}}{n_{2}}}\right\rfloor}\leq\ell^{(2)}\leq{\left\lfloor{\frac{\mathfrak{p}_{1}}{n_{1}}}\right\rfloor}~{},$ (5.35)
where we introduced
$\mathfrak{p}_{1}=t_{1}\mathfrak{m}-{\left({r/2}\right)}~{},$ $\mathfrak{p}_{2}=t_{2}\mathfrak{m}+{\left({r/2}\right)}~{},$
$j^{(1)}=\ell^{(1)}+{\left({r-2}\right)}c_{R}~{},$ $j^{(2)}=\ell^{(2)}+r\,c_{R}~{}.$ (5.36)
The intervals (5.35) are mutually exclusive: if
$-{\left\lfloor{-\frac{\mathfrak{p}_{2}}{n_{2}}}\right\rfloor}\leq{\left\lfloor{\frac{\mathfrak{p}_{1}}{n_{1}}}\right\rfloor}~{},$ (5.37)
then only the modes $\Phi_{m_{1}^{(2)},m_{2}^{(2)}}$ contribute, giving
@@
$Z_{\rm 2d}^{\Phi}$ $=\prod_{\ell^{(2)}=-{\left\lfloor{-\mathfrak{p}_{2}/n_{2}}\right\rfloor}}^{{\left\lfloor{\mathfrak{p}_{1}/n_{1}}\right\rfloor}}{\left({{\omega}_{\rm 2d}\,\ell^{(2)}+\mathfrak{m}+{\rm i}\,\sigma_{\rm 2d}}\right)}^{-1}~{}.$ (5.38)
On the other hand, if
$1+{\left\lfloor{\frac{\mathfrak{p}_{1}}{n_{1}}}\right\rfloor}\leq-1-{\left\lfloor{-\frac{\mathfrak{p}_{2}}{n_{2}}}\right\rfloor}~{},$ (5.39)
then only the eigenfunctions $\mathcal{B}_{m_{1}^{(1)},m_{2}^{(1)}}$ contribute, providing